hen only a resistor is connected across the terminals of an ac generator (113 V) that has a fixed frequency, there is a current of 0.505 A in the resistor. When only an inductor is connected across the terminals of this same generator, there is a current of 0.408 A in the inductor. When both the resistor and the inductor are connected in series between the terminals of this generator, calculate the following.

(a) the impedance of the series combination Ω

(b) the phase angle between the current and the voltage of the generator °
Answer #1


I1 = E/R


Resistance R = E/I1 = 113/0.505 = 223.8 ohm

when only inductor is connected

I2 = E/(XL) = E/(w*L)


inductive reactance = XL = E/I2 = 113/0.408 = 277 ohms

impedance Z = sqrt(R^2+(XL)^2)


Z = sqrt(223.8^2+277^2) = 356.1 ohm <<<<<--------ANSWER


================

part (b)

phase angle = atan(XL/R)


phase angle = atan(277/223.8) = 51 degrees <<<<<--------ANSWER